OpenStudy (anonymous):

All variables in the following problem represent positive real numbers. Simplify each term, and combine, if possible.

OpenStudy (anonymous):

\[5\sqrt{180x^3y^2}-3x \sqrt{125xy^2}\]

OpenStudy (anonymous):

\[5(6xy \sqrt{5x})-3x(5y \sqrt{5x})\]?

OpenStudy (shiraz14):

30xy√(5x) - 15xy√(5x) = 15xy√(5x) (answer)

OpenStudy (anonymous):

\[30xy \sqrt{5x}-15xy \sqrt{5x}\]

OpenStudy (shiraz14):

@cookiibabii93 : Yes

OpenStudy (anonymous):

Ok, so I was doing it right! Thank you !!!

OpenStudy (shiraz14):

@cookiibabii93 : You're welcome, but you need to further evaluate it, to get the answer [i.e. 15xy√(5x)].

OpenStudy (anonymous):

I know, just subtract 30-15

OpenStudy (shiraz14):

@cookiibabii93 : You're correct.